OpenStudy (anonymous):

Given that P = (5, 8) and Q = (6, 9), find the component form and magnitude of PQ

OpenStudy (anonymous):

Note that a vector \(V\) from point \(P\) to \(Q\) is: \[ Q-P=V \]So, we have that: \[ V=Q-P=<6,9>-<5, 8>=<1,1> \]So: \[ V=<1,1>=1{\bf{i}}+1\bf{j} \]The magnitude is given by: \[ |V|=\sqrt{x^2+y^2}=\sqrt{1+1}=\sqrt{2} \]

OpenStudy (anonymous):